40m QIL Cryo_Lab CTN SUS_Lab TCS_Lab OMC_Lab CRIME_Lab FEA ENG_Labs OptContFac Mariner WBEEShop
  40m Log  Not logged in ELOG logo
Entry  Mon Jan 23 15:16:52 2017, gautam, Update, Optical Levers, ETMY Oplev HeNe needs to be replaced ETMY_Oplev.pdf
    Reply  Mon Jan 23 17:24:26 2017, Steve, Update, Optical Levers, ETMY Oplev HeNe is running hot ETMY_oplev__lasers.png
       Reply  Tue Jan 24 17:52:15 2017, Craig, Update, Optical Levers, ETMY Oplev HeNe is replaced 6x
          Reply  Wed Mar 1 08:28:04 2017, Steve, Update, Optical Levers, oplev laser RIN RIN_1103P.pngRIN_HeNe.png
             Reply  Thu Mar 2 13:59:04 2017, Steve, Update, Optical Levers, oplev laser RIN 1103P@2d.png
                Reply  Thu Mar 2 17:58:45 2017, rana, Update, Optical Levers, oplev laser RIN 
                   Reply  Fri Mar 3 17:24:21 2017, Steve, Update, Optical Levers, oplev laser RIN 3march17.pdf
                      Reply  Wed Mar 8 15:28:37 2017, Steve, Update, Optical Levers, oplev laser RIN RIN.jpgRIN_1103P_rotated.pdf
                      Reply  Mon Apr 10 09:49:32 2017, Steve, Update, Optical Levers, oplev laser RIN test planning 
                         Reply  Mon Apr 10 14:21:57 2017, rana, Update, Optical Levers, oplev laser RIN test planning 
       Reply  Tue May 30 13:36:58 2017, Steve, Update, Optical Levers, ETMY Oplev HeNe is replaced oplevETMY120d.png
          Reply  Tue Jan 15 18:16:00 2019, gautam, Update, Optical Levers, ETMY Oplev HeNe needs replacement OLRIN.pdfOLsums.png
             Reply  Fri Jan 18 12:52:07 2019, gautam, Update, Optical Levers, BS/PRM Oplev HeNe replaced OLRIN_20190118.pdf
Message ID: 14404     Entry time: Fri Jan 18 12:52:07 2019     In reply to: 14402
Author: gautam 
Type: Update 
Category: Optical Levers 
Subject: BS/PRM Oplev HeNe replaced 

I replaced the BS/PRM Oplev HeNe with one of the heads from the SP table where Steve was setting up the OL RIN/pointing noise experiment. The old one was dead. The new one outputs 3.2 mW of power, I've labelled it with this number, serial number and date of replacement. The beam comes out of the vacuum chamber for both the BS and PRM, and the RIN spectra (Attachment #1) look alright. The calibration into urad and loop gains possibly have to be tweaked. Since the beam comes out of vacuum, I say that we shouldn't open the BS/PRM chamber for this vent - we don't have a proper plan for the in-air layout yet, so we can add this to the list of to-dos for the next vent.

I think we are down to our last spare HeNe head in the lab - @Chub, please look into ordering some more, the ITMX HeNe is going to need replacement soon.

Attachment 1: OLRIN_20190118.pdf  43 kB  | Hide | Hide all
OLRIN_20190118.pdf
ELOG V3.1.3-